OpenStudy (anonymous):

Please help! Find the remaining parts of the right triangle having side a = 12 and side b = 16. Give the results properly rounded off. B =

OpenStudy (anonymous):

u already gave us what B is?

OpenStudy (anonymous):

may b use the Pythagorean theorem! 12^2 + 16 ^2 = B^2 then 400 = B^2 20 = B

OpenStudy (anonymous):

Hmm... My options are: a. 52 degrees b. 37 degrees c. 49 degrees d. 41 degrees

OpenStudy (anonymous):

oh! you want the angle measurement! i thought you needed the side! ok! 1 sec!

OpenStudy (anonymous):

use the tangent function...tanB = 16/12 then B = inverse tan(1.33) = 53... A seems like the closest one!

OpenStudy (anonymous):

Thanks so much!